OpenStudy (anonymous):

find the values of the trigonometric functions of theta from the information given. cot theta =1/4, sin theta <0. a)sin theta_____ b)cos theta____c)csc theta__ d)sec theta__ e)cot theta__

OpenStudy (anonymous):

E :)

OpenStudy (anonymous):

Use the trig identity: sin^2x = 1/(1 = cot^2 x) sin^2 x = 1/(1 + 1/16) = 16/17 --> sin x = V16/17 = sin 70.25 deg cos^2 x = 1 - (16/17)^2 = 33/289 -> cos x = 0.34 sec x = csc x =